# Site-Decorated Model for Unconventional Frustrated Magnets with Ultranarrow Phase Crossover and Spin Reversal Transition

### v1

@@ III Discussion
The present results have the following immediate impacts and stimulate a series of subsequent work:
@@
I have not solved the critical outstanding problem of the 2D and 3D Ising models in an external magnetic field. I used the salient feature of $h_{\mathrm{eff}}=0$ in the site-decorated Ising models to exactly reveal the spin reversal transition at $T_{0}$ in the 2D and 3D Ising model in an external magnetic field. Furthermore, reexamining the 2D or 3D bond-decorated ferrimagnetic Ising models with ferromagnetic $J$ [Fig. 1(c)], the spin reversal transition occurs, too (not shown), though not as obvious. The present result calls for Monte Carlo simulation and micromagnetic calculation of the complete $h-T$ phase diagram of such decorated 2D and 3D Ising models.
@@
Traditionally, spin reversal in a ferromagnet or ferrimagnet is achieved by applying a magnetic field in the opposite direction. This operation, e.g., in magneto recording, has to overcome the coercive force of the magnetic material. It is often needed to increase the local temperature to so close to the Curie temperature $T_{c}$, typically by laser illumination, that a small applied magnetic field can induce the magnetization reversal [1]. Therefore, the present spin reversal mechanism is attractive as it is accessible by a slight change in temperature or the magnetic field. The transition is driven by the hidden half-ice, half-fire state, in which the sublattice of the decorated, type-b spins provides a macroscopic degeneracy of $2^{N}$, causing the extremely sensitive, energy-efficient response of the type-a spins at $T_{0}$. It is worth mentioning that such response can occur for $h\to 0$, especially in two and three dimensions. The abrupt change in the system’s behavior for $h=0$ and $h\to 0$ is rare and can be used to measure the absolute zero magnetic field.
For the materialization of the site-decorated models, mixed $d$-$f$ compounds are material candidates [1, 15], in which the type-a and type-b spins can be $3d$-$5d$ and $4f$ elements, respectively, as the wave functions of the $3d$-$5d$ electronic orbitals are more likely to overlap than the $4f$ orbitals. The $4f$ magnetic moments normally do not talk with each other and act as the decorated spins. Besides, it is not uncommon to find larger magnetic moments in the $4f$ elements than the $3d$-$5d$ ones to satisfy $\mu_{b}>\mu_{a}$. It is also natural to explore optical lattices [23] and neural networks [24, 25], among others, for they have been built to study other Ising models.
@@
Another way to meet the criterion of $\eta\equiv\mu_{a}/\mu_{b}<1$ is to do multi-side ($M$-side) site decoration. $M=1$ in Figs. 1(b)(d). The examples of $M=2$ and $4$ are illustrated in Fig. 5. Then, the effective magnetic field becomes
$h_{\mathrm{eff}}=h+\frac{M}{2\beta\mu_{a}}\left(\ln a-\ln b\right).$ (11)
@@
This results in the replacement of $\mu_{a}$ by $\mu_{a}/M$ in Equations (7) and (8), i.e., $\eta$ is effectively reduced by a factor of $M$. Consequently, $T_{0}$ could be increased dramatically.
Last but not least, the fact that site and bond decorations are geometrically compatible would significantly enlarge the possibility of functional materials design. The advancement in our knowledge of the complexity and vast possibility of the Ising model, a basic model for cooperative phenomena in various physical, biological, economical, and social systems, by the way, provides a fundamental playground for machine learning and AI for sciences.

### v2

@@ III Discussion
The present results have immediate implications and inspire a series of subsequent studies, as outlined below.
@@
I have not yet solved the long-standing problem of the 2D and 3D Ising models in a magnetic field. However, by leveraging the key feature of $h_{\mathrm{eff}}=0$ in the site-decorated Ising models, I have exactly revealed the spin reversal transition at $T_{0}$ in the 2D and 3D Ising models under an external field. Furthermore, a reexamination of 2D and 3D bond-decorated models with ferromagnetic $J$ [Fig. 1(c)] also reveals a spin reversal transition (not shown), though it is less obvious. These findings highlight the need for Monte Carlo simulations [27] to map the complete $h-T$ phase diagram of such decorated 2D and 3D Ising models.
@@
Traditionally, spin reversal in ferromagnets and ferrimagnets is induced by applying a magnetic field in the opposite direction. In applications such as magnetic recording, overcoming the coercive force often requires increasing the local temperature close to the Curie temperature $T_{c}$, typically through laser illumination, so that a small applied field can trigger magnetization reversal [2]. By contrast, the spin reversal mechanism presented here is particularly attractive, as it can be achieved through a slight change in temperature or magnetic field. This transition is driven by the hidden half-ice, half-fire state, in which the decorated type-b sublattice provides a macroscopic degeneracy of $2^{N}$, resulting in an extremely sensitive and energy-efficient response of the type-a spins at $T_{0}$. Notably, this response persists as $h\to 0$, particularly in two and three dimensions. The abrupt contrast in system behavior between $h=0$ and $h\to 0$ is rare and could serve as a method for measuring absolute zero magnetic field.
For the experimental realization of site-decorated models, mixed $d$-$f$ or $3d$-$5d$ compounds are promising candidates [2, 19]. In these systems, the type-a and type-b spins correspond to $d$ and $4f$ elements, respectively, since the electronic wavefunctions of $d$ orbitals exhibit greater overlap than those of $4f$ orbitals. The weak interaction between $4f$ moments allows them to act as decorated spins, and the typically larger magnetic moments of $4f$ elements compared to $d$ ones naturally satisfy the condition $\mu_{b}>\mu_{a}$. A similar argument applies to $3d$-$5d$ compounds. Additionally, optical lattices [28] and neural networks [29, 30], both of which have been employed to study other Ising models, present alternative platforms for exploring site-decorated systems.
@@
Another way to satisfy the condition $\mu_{a}/\mu_{b}<1$ is through multi-side ($M$-side) site decoration. Figs. 1(b)(d) represent $M=1$. The examples of $M=2$ and $4$ are illustrated in Fig. 5. Then, the effective magnetic field becomes
$h_{\mathrm{eff}}=h+\frac{M}{2\beta\mu_{a}}\left(\ln\boxed{+}-\ln\boxed{-}\right).$ (11)
@@
This results in the replacement of $\mu_{a}$ by $\mu_{a}/M$ in Equations (7) and (8), i.e., $\mu_{a}/\mu_{b}$ is effectively reduced by a factor of $M$. Consequently, $T_{0}$ can increase dramatically.
Finally, the geometric compatibility of site and bond decorations significantly expands the potential for functional materials design. Moreover, advancing our understanding of the complexity and vast possibilities of the Ising model—a fundamental framework for cooperative phenomena across physical, biological, economic, and social systems—provides a crucial foundation for machine learning and AI applications in science.
IV Summary
In summary, site decoration has been introduced as a novel scheme for frustrated magnets, enabling exotic phase-switching phenomena in any dimension and yielding a rare exact result for 2D and 3D systems in a magnetic field. These discoveries establish a new paradigm for designing materials and devices with extreme sensitivity, where a hidden frustrated state—the half-ice, half-fire state—drives an abrupt spin reversal even under an infinitesimal applied field.